Vicky Le, an employee of Sweet Shoppe Industries, receives a bonus of $3,800 for her stellar work. Her boss wants Vicky to receive $3,800 on the check. She contributes 3% of her pay in a pre-tax deduction to her 401(k). Calculate the gross pay amount that would result in $3,800 paid to Vicky. (The tax rate on bonuses is 28%. The Social Security (6.2%) and Medicare taxes (1.45%) must be added to this rate.)

Gross Pay ________
Answer #1

Total rate of deductions from gross pay = 3% + 28%+6.2%+1.45% = 38.65%

Net pay as a ratio of gross pay = 1- 38.65% = 61.35%

Net pay required = 3,800

Gross pay = 3,800/ 61.35% = 6,193.969